OpenStudy (anonymous):

by direct integration find the area of that portion of the sphere x^2+y^2+z^2=25 which falls between planes z=3 and z=4

OpenStudy (turingtest):

the region D that we will be integrating over is\[x^2+y^2+3^2=25\implies x^2+y^2=16\]

OpenStudy (turingtest):

I suppose we can put this into cylindrical coordinates

OpenStudy (anonymous):

how?

OpenStudy (turingtest):

\[\iiint dV=\iint\limits_D f(x,y)dA=\iint\limits_D 25-x^2-y^2dA\]the region as I mentioned is a circle of radius 4\[D=\{(r,\theta)|0\le r\le 4;0\le\theta\le2\pi\}\]so we get\[\iint\limits_D 25-r^2dA=\int_0^{2\pi}\int_0^4 25r-r^3drd\theta\]

OpenStudy (anonymous):

no radius is between 4 and 3

OpenStudy (turingtest):

no, I put the region we are integrating over into polar coordinates

OpenStudy (turingtest):

the radius is a constant of taken to be from the origin, because this is a sphere

OpenStudy (turingtest):

if taken*

OpenStudy (turingtest):

the question hinges on what coordinate system we are using

OpenStudy (anonymous):

yes, when we changed to polar coordinate, then the radius will from 4 to 3 because for z=3 it will give us radius of 4 and for z=4 then it gives radius of 3

OpenStudy (turingtest):

I rewrote z as a function of x and y and put the region of integration into polar coordinates z is a function of x and y (i.e. where we are in the x,y plane) so son't confuse that with radius, which we only used for the region

OpenStudy (anonymous):

okay

OpenStudy (turingtest):

the region is the whole circle because we have to chose a shape that lies *below* the thing we are integrating (like the shadow that would be made by shining a light on the object toward the xy-plane)

OpenStudy (anonymous):

thanks

OpenStudy (turingtest):

I was also able to simplify the integral by plugging\[x=r\cos\theta\]\[y=r\sin\theta\]so as you can see radius is again a function of x and y

OpenStudy (turingtest):

welcome